I need help with this 1 question.
For events A and B, P(A) = 4/3 and P(B) = 1/4. Also, P(A and B) = 1/5. Are A and B independent? Explain.

Answer by edjones(8007)

P(A|B)=P(A and B)/P(B)= 1/5 / 1/4 = 4/5
P(A|B)<> P(A) so it is not independent.
If it were independent the probability of event A would be the same regardless of whether event B has occurred.
4/5<>4/3
.
Ed